OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

m=3 x1=1 y1=-2

OpenStudy (anonymous):

substitute in the numbers

OpenStudy (mokeira):



OpenStudy (anonymous):

Write an equation in slope-intercept form for the line shown.

OpenStudy (anonymous):

NO idea how to do the equation with graph

OpenStudy (anonymous):

what is the y-intercept?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

yes now do you know how to find slope on a graph?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

My course doesn't explain it very well, it's frustrating me.

OpenStudy (mokeira):

lol....you have been given two points, now just find the slope can you continue from there?

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\] you're given 2 coordinates

OpenStudy (mokeira):

\[slope=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

2,-1 0,-4 x1=2 y1=-1 x2=0 y2=-4

OpenStudy (anonymous):

y = 3/2 x - 4?

OpenStudy (anonymous):

\[\frac{ -4+1 }{ 0-2 }\] yup :)

OpenStudy (anonymous):

Sweet, thank you both, learning more in 5 minutes from here than my whole section that is attempting to teach this.

OpenStudy (anonymous):

wait until you get into calculus :) thatll be fun

OpenStudy (anonymous):

Not looking forward to it D:

OpenStudy (anonymous):

It's semi easy quadratic formula took me 2 hours to understand it but im also still in algebra